Circular No. 7272

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ee IN IC 5179 J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by M. Wischnjewsky, on a 20-min unfiltered T-Max 400 film taken by L. Gonzalez on Oct. 7.15 UT with the University of Chile’s Maksutov telescope, of a supernova (m_{pv} ∼ 17.5) located at α = 22^{h}16^{m}09^{s}.40, δ = −36^{o}50′31′′.5 (equinox 2000.0), which is ∼ 10′′ east and 10′′ south of the nucleus of IC 5179. The supernova was confirmed by P. Candia from B,V,I images taken with the Cerro Tololo 0.9-m telescope on Oct. 8.12. Maza obtained a spectrum (range 320–800 nm; resolution 0.4 nm) of SN 1999ee on Oct. 9.10 using the 4-m Blanco telescope at Cerro Tololo, which reveals that the supernova is a type-Ia event with the characteristic Si II 635.5-nm line having an expansion velocity of 15 700 km/s. The spectrum has a blue continuum, indicating the supernova to be several days before maximum. Hamuy and C. Lidman obtained a spectrum (range 1000–2200 nm; resolution ∼ 500, S/N ∼ 7) on Oct. 9.10, at the European Southern Observatory with the Very Large Telescope (+ ISAAC), which shows a featureless continuum. At the redshift of the host galaxy (3447 km/s), a type-Ia supernova could reach a peak V magnitude near = 14.5.

NOVAE IN M31 A. V. Filippenko, R. T. Chornock, A. L. Coil, D. C. Leonard, and W.

D. Li, University of California at Berkeley, write that CCD spectra (range 420–700 nm) obtained on Sept. 10 and 17 UT with the 3-m Shane reflector at Lick Observatory confirm the novae in M31 reported on IAUC 7218 and 7249, respectively. Prominent hydrogen Balmer emission lines are visible, with weaker Fe II emission.

Li, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) also reports the discovery, with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on unfiltered images taken on Oct. 3.2 and 4.2 UT, of another nova (mag ∼ 17.5) in M31. An image taken on Oct. 2.2 also showed a hint of the nova. The magnitude estimate (∼ 17.5) is rough because of the nova’s proximity to the bright nucleus of M31. The new object is located at α = 0^{h}42^{m}49^{s}.7, δ = +41^{o}16′32′′ (equinox 2000.0), which is ∼ 61′′ east and 24′′ north of the nucleus of M31. A KAIT image of M31 on Oct. 1.2 (limiting mag ∼ 19.0) showed nothing at the position of the nova. Filippenko and Chornock report that a noisy CCD spectrum (range 420–700 nm), obtained on Oct. 8 with the 3-m Shane reflector, confirms the nova; hydrogen Balmer emission lines are present.

1999 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green